# Large Language Model Inference Acceleration: A Comprehensive Hardware Perspective

### v2

@@ 3.2.4 FPGA
Weight Sparsity. FlightLLM [120] is the first real FPGA-based LLM accelerator which proposes a configurable sparse DSP chain to support different sparsity patterns with high computation efficiency. Then, it proposes an always-on-chip decode scheme to boost memory bandwidth with mixed-precision support. Finally, it proposes a length adaptive compilation method to reduce the compilation overhead. For Llama2-7B model, FlightLLM achieves 55 tokens/s and 92.5 tokens/s with batch size 1 on the Xilinx Alveo U280 FPGA and Versal VHK158 FPGA, respectively. EdgeLLM [121] integrates 4-bit weight-only quantization and utilizes log-scale structural sparsity for weight parameters in the matrix multiplication operator. For ChatGLM2-6B model, it achieves average 67 tokens/s with 55.07W power consumption on AMD Xilinx VCU128 FPGA.
3.2.5 ASIC
@@
Attention Sparsity. Spatten [148] leverages token sparsity, head sparsity, and quantization opportunities to reduce the attention computation and memory access. It assesses the cumulative importance of each word by aggregating the attention matrix columns, subsequently pruning tokens with minimal cumulative significance from the input in subsequent layers. It provides 2.88 TOPS computing power and 8.3W power consumption by considering all memory access under 40nm. Experimental results shows that for GPT2-Medium, it can achieve about 35.86 tokens/s in decode stage. TF-MVP [149] quantitatively analyzes sparsity patterns of pruned-transformer models with the cutting-edge fine-grained pruning scheme for the first time and presents the mixed-length vector pruning (MVP) procedure by utilizing this direction strength. From hardware perspective, it introduces the TF-MVP architecture, a sparsity-aware cost-efficient accelerator design dedicated to the proposed pruned-transformer models. Implemented in a 28nm CMOS technology at 400MHz, TF-MVP provides 0.835 TOPS with 1.721W on-chip power consumption for accelerating GPT-2 small model. SOFA [150] predicts attention sparsity by using log-based add-only operations to avoid the significant overhead of prediction. Then, a distributed sorting and a sorted updating FlashAttention mechanism are proposed with a cross-stage coordinated tiling principle, which enables fine-grained and lightweight coordination among stages, helping optimize memory access and latency. SOFA provides 24.423 TOPS computing power and 3.4W power consumption under 28nm. For Llama2-7B, compared to NVIDIA A100 GPU, it achieves 9.5$\times$ inference speedup in prefill stage.
3.2.6 PIM/NDP
@@
Weight Sparsity. LauWS [151] proposes an unstructured sparsity method for NDP systems and is evaluated on a practical GDDR6-based bank NDP. Not only the overall features of the total matrix but the features of the local region are preserved by ignoring non-feature values as much as possible, which is beneficial to the trade-off between high sparsity and least accuracy loss. Compared to dense models, it achieves 1.23$\times$ and 1.24$\times$ speedup for GPT-2 small model and OPT-125M at 50% sparsity, respectively. Sharda et al. [129] propose to use the capacitorless 3D stackable DRAM to store much larger LLMs compared to conventional DRAM at higher density. To reduce the intermediate data size, they propose to use a layer-wise sparsity-quantization hybrid (LSQH) algorithm, which induces sparsity based on calculations performed using low-bit quantization to reduce both the energy consumption and the data storage requirements. Finally, a 3D heterogeneously integrated accelerator is designed by stacking a 3D DRAM with logic dies designed in the 3nm technology node at 1GHz. The evaluation shows that for Llama2-13B, it achieves 163k tokens/s in prefill stage with 193W power consumption.
@@
Attention Sparsity. HARDSEA [152] proposes an attention sparsity method by predicting lightweight token relevance and design a hybrid analog-ReRAM and digital-SRAM in-memory computing accelerator. It employs ReRAM-CIM, whose precision is sensitive to circuit non-idealities, to take charge of token relevance prediction where only computing monotonicity is demanded. The SRAM-CIM, utilized for exact sparse attention computing, is reorganized as an on-memory-boundary computing scheme, thus adapting to irregular sparsity patterns. Experimental results show that HARDSEA prunes BERT and GPT-2 small model to 20% sparsity without accuracy loss, achieving 5.8$\times$–6.7$\times$ speedup over NVIDIA RTX 3090 GPU.
@@
Figure 6: LLM ($\sim$ 7 billion parameters) decode stage throughput (batch size 1) vs power on different platforms with sparsity.
3.2.7 Comparison
@@
In Figure 6, for sparsity, power consumption ranges from 30W to 400W, with inference speeds between 8.67 tokens/s and 92.5 tokens/s. The energy efficiency ranges from 0.069 tokens/J to 1.421 tokens/J. TF-MVP [149] (ASIC) achieves the lowest power consumption with 30.0167W and FlightLLM [120] (FPGA) achieves the highest throughput with batch size 1 (though it also applies quantization). And TF-MVP (ASIC) achieves the highest energy efficiency with 1.421 tokens/J.
@@
3.3 Fast Decoding
3.3.1 Overview
Traditional autoregressive decoding typically generates text token by token, choosing only the highest probability token at each step, known as greedy sampling. While this approach is simple and easy to implement, it may lead to a lack of diversity and creativity in the generated results. Another autoregressive method called nucleus sampling (or top-p sampling) [218] considers multiple candidates during generation by setting a cumulative probability threshold p, allowing for sampling within a certain range. Although this method offers more diversity than greedy sampling, it still operates in a step-by-step generation manner. Currently, fast decoding techniques can be mainly divided into two categories: speculative decoding and skip layer.
@@ 3.3.1 Overview
Figure 7: Fast decoding.
Table 5: Fast decoding on CPU, GPU, FPGA, ASIC, and PIM/NDP
@@ 3.3.2 GPU
Speculative Decoding. Speculative decoding [154] is proposed to overcome the inherently sequential process in the autoregressive decoding of LLM. The essential decoding mechanism is to make predictions (i.e., draft tokens) parallelly for multiple time steps and then select the longest prefix verified by a scoring model as the final output. Lookahead decoding [155] adopts the Guess-and-Verify paradigm as the whole decoding mechanism which generates the draft tokens by $n$-gram method and verifies the draft tokens during the forward at the same time. For Llama-7B and Llama-13B models with different datasets, it achieves 65.12-94.51 tokens/s and 56.01-90.05 tokens/s on a single NVIDIA A100-80GB GPU. To improve the acceptance rate while maintaining generation quality, many works focus on the scheme of generating draft tokens. Medusa [156] adds extra decoding heads to generate multiple subsequent tokens in parallel and uses a tree-based attention mechanism to construct multiple candidate continuations and verify them simultaneously in each decoding step. For Vicuna-7B and Vicuna-13B, it achieves 129.86 tokens/s and 98.54 tokens/s on a single NVIDIA A100-80GB GPU. EAGLE [157, 158] selects a single transformer layer with the same configuration as LLM as the draft model to make predictions autoregressively and combines the feature and token embedding as the input of the draft model. During the verification phase on the target model, EAGLE chooses the tree-based attention mechanism similar to Lookahead and Medusa to ensure the correct relationship between the draft tokens. For Vicuna-7B, Vicuna-13B, Llama2-Chat-7B, and Llama-2-Chat-13B models, it achieves 139.95 tokens/s, 132.31 tokens/s, 133.98 tokens/s, and 156.80 tokens/s on a single NVIDIA A100-80GB GPU. Based on the basic paradigm of speculative decoding for prediction by draft models and verification by target models, some studies explore the optimizations on the modules in it. Ouroboros [159] constructs a phrase candidate pool from the verification process of LLMs to provide candidates for the draft token generation of the draft model. Different from Medusa and EAGLE, Ouroboros uses the smaller LLMs (e.g., DeepSeek-7B) as the draft models for the target models (e.g., DeepSeek-34B). For Yi-34B, DeepSeek-34B and CodeLlama-34B models, it achieves 61.20 tokens/s, 41.00 tokens/s and 39.2 tokens/s on a single NVIDIA A100-80GB GPU. During the prediction phase, Sequoia [160] introduces a dynamic programming algorithm and a hardware-aware tree optimizer to find the optimal tree structure based on the runtime features and the given hardware platform. During the verification phase, Sequoia uses a novel sampling and verification method that outperforms prior work across different decoding temperatures. For Llama2-7B and Llama2-13B models, it achieves 169.68 tokens/s and 149.20 tokens/s on a single NVIDIA A100-80GB GPU.
@@
The studies mentioned above all require the help of an auxiliary model (e.g., a single transformer layer in EAGLE, several Medusa heads in Medusa, or a smaller LLM like Llama2-7B in Ouroboros) or the statistical methods (e.g., $n$-gram in Lookahead) to generate the predicted tokens and then the target model is utilized to verify the predicted tokens. Some other studies [161, 162, 163] also explore the LLM inference acceleration without the need of auxiliary models, called self-speculative decoding. Draft&Verify [161] generates draft tokens by selectively skipping certain intermediate layers of LLMs. Subsequently, the draft tokens will be verified in one forward pass. For the Llama2-13B model, it achieves 62.23 tokens/s on a single NVIDIA A100-40GB GPU. Kangaroo [162] adopts a fixed shallow sub-network of the LLM as the draft model, with the remaining layers serving as the target model. To enhance the representation ability of the draft model (i.e., the shallow sub-network), it trains an adapter module to follow the sub-network. For Vicuna-7B and Vicuna-13B models, it achieves 138.14 tokens/s and 105.89 tokens/s on a single NVIDIA A100-80GB GPU. LayerSkip [163] proposes to exit at early layers and verify and correct with remaining layers of the model. During training, it applies layer dropout with low dropout rates for earlier layers and higher dropout rates for later layers, and adds an early exit loss to increase the accuracy of early exit at earlier layers. For the Llama2-13B model, it achieves 66.37 tokens/s on a single NVIDIA H100 GPU. LLMA [153] is motivated by the observation that there are abundant identical text spans between the decoding result by an LLM and the reference that is available in many real-world scenarios (e.g., retrieved documents). LLMA first selects a text span from the reference and copies its tokens to the decoder and then efficiently checks the tokens’ appropriateness as the decoding result in parallel within one decoding step. For Llama-7B and Llama-13B models, it achieves 59.2 tokens/s and 41.1 tokens/s on a single NVIDIA V100 GPU, respectively.
@@
Skip Layer. The skip layer method [164, 165, 166] is proposed based on the idea that not all layers of LLMs are necessary during inference. AdaInfer [164] statistically analyzes the activated layers across tasks and proposes a simple algorithm to determine the inference termination moment based on the input instance adaptively. Due to the introduction of the overhead in each layer, which is not friendly to the decoding phase, AdaInfer only takes some Benchmarks for the Q&A tasks and achieves 25.2 tokens/s for Llama2-7B on a single NVIDIA V100 GPU. Based on the basic dataflow of Adainfer, RAEE [165] proposes to build the retrieval database to store the token information offline and leverages the information of the retrieved similar token by searching the pre-built retrieval database to guide the backbone model to exit at the layer. For the LM-BFF which is finetuned based on RoBERT-large-350M, RAEE achieves 26.23 tokens/s. The studies mentioned above both use continuous shallow layers for inference, called early exiting. MOD [166] decides whether to skip the current layer or not by pretraining the model to add a router in each layer like Mixture-of-Experts. This achieves a dynamic selection of partial layers for computation instead of forcing continuous layers. The model of MOD is not open source and no corresponding results on throughput are given in the paper.
3.3.3 ASIC
@@
Speculative Decoding. C-Transformer [167] adopts a big-little network, which is composed of the original GPT-2 big model and a 1/10$\times$ smaller model, and a reconfigurable homogeneous architecture to increase hardware utilization and energy efficiency. During inference, only the little model computation is performed, and if the prediction probability of a specific token is over a predefined threshold, the big model computation is skipped leading to memory access reduction. Then, workloads are divided into two domains: adder-based spike domain which is efficient for small input values, and multiplier-based non-spike domain which is efficient for large input values. It has 1.431W power consumption and is fabricated in 28nm CMOS technology. For GPT-2 large model, C-Transformer achieves 2146.75 tokens/s in prefill stage.
@@
Figure 8: LLM ($\sim$ 7 billion parameters) decode stage throughput (batch size 1) vs power on different platforms with fast decoding.
3.3.4 PIM/NDP

### v3

@@ 3.2.4. FPGA
Weight Sparsity. FlightLLM (zeng2024flightllm, ) is the first real FPGA-based LLM accelerator which proposes a configurable sparse DSP chain to support different sparsity patterns with high computation efficiency. Then, it proposes an always-on-chip decode scheme to boost memory bandwidth with mixed-precision support. Finally, it proposes a length adaptive compilation method to reduce the compilation overhead. For Llama2-7B model, FlightLLM achieves 55 tokens/s and 92.5 tokens/s with batch size 1 on the Xilinx Alveo U280 FPGA and Versal VHK158 FPGA, respectively. EdgeLLM (huang2024edgellm, ) integrates 4-bit weight-only quantization and utilizes log-scale structural sparsity for weight parameters in the matrix multiplication operator. For ChatGLM2-6B model, it achieves average 67 tokens/s with 55.07W power consumption on AMD Xilinx VCU128 FPGA.
3.2.5. ASIC
@@
Attention Sparsity. Spatten (wang2021spatten, ) leverages token sparsity, head sparsity, and quantization opportunities to reduce the attention computation and memory access. It assesses the cumulative importance of each word by aggregating the attention matrix columns, subsequently pruning tokens with minimal cumulative significance from the input in subsequent layers. It provides 2.88 TOPS computing power and 8.3W power consumption by considering all memory access under 40nm. Experimental results shows that for GPT2-Medium, it can achieve about 35.86 tokens/s in decode stage. TF-MVP (yoo2023tf, ) quantitatively analyzes sparsity patterns of pruned-transformer models with the cutting-edge fine-grained pruning scheme for the first time and presents the mixed-length vector pruning (MVP) procedure by utilizing this direction strength. From hardware perspective, it introduces the TF-MVP architecture, a sparsity-aware cost-efficient accelerator design dedicated to the proposed pruned-transformer models. Implemented in a 28nm CMOS technology at 400MHz, TF-MVP provides 0.835 TOPS with 1.721W on-chip power consumption for accelerating GPT-2 small model. SOFA (wang2024sofa, ) predicts attention sparsity by using log-based add-only operations to avoid the significant overhead of prediction. Then, a distributed sorting and a sorted updating FlashAttention mechanism are proposed with a cross-stage coordinated tiling principle, which enables fine-grained and lightweight coordination among stages, helping optimize memory access and latency. SOFA provides 24.423 TOPS computing power and 3.4W power consumption under 28nm. For Llama2-7B, compared to NVIDIA A100 GPU, it achieves 9.5$\times$ inference speedup in prefill stage.
3.2.6. PIM/NDP
@@
Weight Sparsity. LauWS (li2024lauws, ) proposes an unstructured sparsity method for NDP systems and is evaluated on a practical GDDR6-based bank NDP. Not only the overall features of the total matrix but the features of the local region are preserved by ignoring non-feature values as much as possible, which is beneficial to the trade-off between high sparsity and least accuracy loss. Compared to dense models, it achieves 1.23$\times$ and 1.24$\times$ speedup for GPT-2 small model and OPT-125M at 50% sparsity, respectively. Sharda et al. (sharda2024accelerator, ) propose to use the capacitorless 3D stackable DRAM to store much larger LLMs compared to conventional DRAM at higher density. To reduce the intermediate data size, they propose to use a layer-wise sparsity-quantization hybrid (LSQH) algorithm, which induces sparsity based on calculations performed using low-bit quantization to reduce both the energy consumption and the data storage requirements. Finally, a 3D heterogeneously integrated accelerator is designed by stacking a 3D DRAM with logic dies designed in the 3nm technology node at 1GHz. The evaluation shows that for Llama2-13B, it achieves 163k tokens/s in prefill stage with 193W power consumption.
@@
Attention Sparsity. HARDSEA (liu2023hardsea, ) proposes an attention sparsity method by predicting lightweight token relevance and design a hybrid analog-ReRAM and digital-SRAM in-memory computing accelerator. It employs ReRAM-CIM, whose precision is sensitive to circuit non-idealities, to take charge of token relevance prediction where only computing monotonicity is demanded. The SRAM-CIM, utilized for exact sparse attention computing, is reorganized as an on-memory-boundary computing scheme, thus adapting to irregular sparsity patterns. Experimental results show that HARDSEA prunes BERT and GPT-2 small model to 20% sparsity without accuracy loss, achieving 5.8$\times$–6.7$\times$ speedup over NVIDIA RTX 3090 GPU.
@@
Figure 6. LLM ($\sim$ 7 billion parameters) decode stage throughput (batch size 1) vs power on different platforms with sparsity.
3.2.7. Comparison
@@
In Figure 6, for sparsity, power consumption ranges from 30W to 400W, with inference speeds between 8.67 tokens/s and 92.5 tokens/s. The energy efficiency ranges from 0.069 tokens/J to 1.421 tokens/J. TF-MVP (yoo2023tf, ) (ASIC) achieves the lowest power consumption with 30.0167W and FlightLLM (zeng2024flightllm, ) (FPGA) achieves the highest throughput with batch size 1 (though it also applies quantization). And TF-MVP (ASIC) achieves the highest energy efficiency with 1.421 tokens/J.
@@
3.3. Fast Decoding
3.3.1. Overview
Traditional autoregressive decoding typically generates text token by token, choosing only the highest probability token at each step, known as greedy sampling. While this approach is simple and easy to implement, it may lead to a lack of diversity and creativity in the generated results. Another autoregressive method called nucleus sampling (or top-p sampling) (holtzman2019curious, ) considers multiple candidates during generation by setting a cumulative probability threshold p, allowing for sampling within a certain range. Although this method offers more diversity than greedy sampling, it still operates in a step-by-step generation manner. Currently, fast decoding techniques can be mainly divided into two categories: speculative decoding and skip layer.
@@ 3.3.1. Overview
Figure 7. Fast decoding.
Table 5. Fast decoding on CPU, GPU, FPGA, ASIC, and PIM/NDP
@@ 3.3.2. GPU
Speculative Decoding. Speculative decoding (stern2018blockwise, ) is proposed to overcome the inherently sequential process in the autoregressive decoding of LLM. The essential decoding mechanism is to make predictions (i.e., draft tokens) parallelly for multiple time steps and then select the longest prefix verified by a scoring model as the final output. Lookahead decoding (lookahead, ) adopts the Guess-and-Verify paradigm as the whole decoding mechanism which generates the draft tokens by $n$-gram method and verifies the draft tokens during the forward at the same time. For Llama-7B and Llama-13B models with different datasets, it achieves 65.12-94.51 tokens/s and 56.01-90.05 tokens/s on a single NVIDIA A100-80GB GPU. To improve the acceptance rate while maintaining generation quality, many works focus on the scheme of generating draft tokens. Medusa (medusa, ) adds extra decoding heads to generate multiple subsequent tokens in parallel and uses a tree-based attention mechanism to construct multiple candidate continuations and verify them simultaneously in each decoding step. For Vicuna-7B and Vicuna-13B, it achieves 129.86 tokens/s and 98.54 tokens/s on a single NVIDIA A100-80GB GPU. EAGLE (eagle1, ; eagle2, ) selects a single transformer layer with the same configuration as LLM as the draft model to make predictions autoregressively and combines the feature and token embedding as the input of the draft model. During the verification phase on the target model, EAGLE chooses the tree-based attention mechanism similar to Lookahead and Medusa to ensure the correct relationship between the draft tokens. For Vicuna-7B, Vicuna-13B, Llama2-Chat-7B, and Llama-2-Chat-13B models, it achieves 139.95 tokens/s, 132.31 tokens/s, 133.98 tokens/s, and 156.80 tokens/s on a single NVIDIA A100-80GB GPU. Based on the basic paradigm of speculative decoding for prediction by draft models and verification by target models, some studies explore the optimizations on the modules in it. Ouroboros (zhao2024ouroboros, ) constructs a phrase candidate pool from the verification process of LLMs to provide candidates for the draft token generation of the draft model. Different from Medusa and EAGLE, Ouroboros uses the smaller LLMs (e.g., DeepSeek-7B) as the draft models for the target models (e.g., DeepSeek-34B). For Yi-34B, DeepSeek-34B and CodeLlama-34B models, it achieves 61.20 tokens/s, 41.00 tokens/s and 39.2 tokens/s on a single NVIDIA A100-80GB GPU. During the prediction phase, Sequoia (chen2024sequoia, ) introduces a dynamic programming algorithm and a hardware-aware tree optimizer to find the optimal tree structure based on the runtime features and the given hardware platform. During the verification phase, Sequoia uses a novel sampling and verification method that outperforms prior work across different decoding temperatures. For Llama2-7B and Llama2-13B models, it achieves 169.68 tokens/s and 149.20 tokens/s on a single NVIDIA A100-80GB GPU.
@@
The studies mentioned above all require the help of an auxiliary model (e.g., a single transformer layer in EAGLE, several Medusa heads in Medusa, or a smaller LLM like Llama2-7B in Ouroboros) or the statistical methods (e.g., $n$-gram in Lookahead) to generate the predicted tokens and then the target model is utilized to verify the predicted tokens. Some other studies (zhang2023draft, ; liu2024kangaroo, ; layerskip, ) also explore the LLM inference acceleration without the need of auxiliary models, called self-speculative decoding. Draft&Verify (zhang2023draft, ) generates draft tokens by selectively skipping certain intermediate layers of LLMs. Subsequently, the draft tokens will be verified in one forward pass. For the Llama2-13B model, it achieves 62.23 tokens/s on a single NVIDIA A100-40GB GPU. Kangaroo (liu2024kangaroo, ) adopts a fixed shallow sub-network of the LLM as the draft model, with the remaining layers serving as the target model. To enhance the representation ability of the draft model (i.e., the shallow sub-network), it trains an adapter module to follow the sub-network. For Vicuna-7B and Vicuna-13B models, it achieves 138.14 tokens/s and 105.89 tokens/s on a single NVIDIA A100-80GB GPU. LayerSkip (layerskip, ) proposes to exit at early layers and verify and correct with remaining layers of the model. During training, it applies layer dropout with low dropout rates for earlier layers and higher dropout rates for later layers, and adds an early exit loss to increase the accuracy of early exit at earlier layers. For the Llama2-13B model, it achieves 66.37 tokens/s on a single NVIDIA H100 GPU. LLMA (yang2023inference, ) is motivated by the observation that there are abundant identical text spans between the decoding result by an LLM and the reference that is available in many real-world scenarios (e.g., retrieved documents). LLMA first selects a text span from the reference and copies its tokens to the decoder and then efficiently checks the tokens’ appropriateness as the decoding result in parallel within one decoding step. For Llama-7B and Llama-13B models, it achieves 59.2 tokens/s and 41.1 tokens/s on a single NVIDIA V100 GPU, respectively.
@@
Skip Layer. The skip layer method (adainfer, ; RAEE, ; MOD, ) is proposed based on the idea that not all layers of LLMs are necessary during inference. AdaInfer (adainfer, ) statistically analyzes the activated layers across tasks and proposes a simple algorithm to determine the inference termination moment based on the input instance adaptively. Due to the introduction of the overhead in each layer, which is not friendly to the decoding phase, AdaInfer only takes some Benchmarks for the Q&A tasks and achieves 25.2 tokens/s for Llama2-7B on a single NVIDIA V100 GPU. Based on the basic dataflow of Adainfer, RAEE (RAEE, ) proposes to build the retrieval database to store the token information offline and leverages the information of the retrieved similar token by searching the pre-built retrieval database to guide the backbone model to exit at the layer. For the LM-BFF which is finetuned based on RoBERT-large-350M, RAEE achieves 26.23 tokens/s. The studies mentioned above both use continuous shallow layers for inference, called early exiting. MOD (MOD, ) decides whether to skip the current layer or not by pretraining the model to add a router in each layer like Mixture-of-Experts. This achieves a dynamic selection of partial layers for computation instead of forcing continuous layers. The model of MOD is not open source and no corresponding results on throughput are given in the paper.
3.3.3. ASIC
@@
Speculative Decoding. C-Transformer (kim202420, ) adopts a big-little network, which is composed of the original GPT-2 big model and a 1/10$\times$ smaller model, and a reconfigurable homogeneous architecture to increase hardware utilization and energy efficiency. During inference, only the little model computation is performed, and if the prediction probability of a specific token is over a predefined threshold, the big model computation is skipped leading to memory access reduction. Then, workloads are divided into two domains: adder-based spike domain which is efficient for small input values, and multiplier-based non-spike domain which is efficient for large input values. It has 1.431W power consumption and is fabricated in 28nm CMOS technology. For GPT-2 large model, C-Transformer achieves 2146.75 tokens/s in prefill stage.
@@
Figure 8. LLM ($\sim$ 7 billion parameters) decode stage throughput (batch size 1) vs power on different platforms with fast decoding.
3.3.4. PIM/NDP